### v4

@@ Abstract
Reasoning over sports videos for question answering is an important task with numerous applications, such as player training and information retrieval. However, this task has not been explored due to the lack of relevant datasets and the challenging nature it presents. Most datasets for video question answering (VideoQA) focus mainly on general and coarse-grained understanding of daily-life videos, which is not applicable to sports scenarios requiring professional action understanding and fine-grained motion analysis. In this paper, we introduce the first dataset, named Sports-QA, specifically designed for the sports VideoQA task. The Sports-QA dataset includes various types of questions, such as descriptions, chronologies, causalities, and counterfactual conditions, covering multiple sports. Furthermore, to address the characteristics of the sports VideoQA task, we propose a new Auto-Focus Transformer (AFT) capable of automatically focusing on particular scales of temporal information for question answering. We conduct extensive experiments on Sports-QA, including baseline studies and the evaluation of different methods. The results demonstrate that our AFT achieves state-of-the-art performance.
@@ 1 Introduction
Sports video analysis has been attracting increasing attention in recent years [45, 19, 17, 49, 26, 38, 22]. While research progress has been made on tasks such as sports action recognition [18, 49, 29], reasoning over sports videos for question answering has not been explored. As humans, we can not only recognize the actions of the players in sports videos but also understand the effects of the players’ actions, explain why a team loses the score, and imagine what would happen under counterfactual situations. Our impressive capabilities in reasoning allow us to answer complex questions related to sports videos, which is crucial in applications such as obtaining crucial statistics in matches for player/team performance qualification, analyzing players’ actions and team strategies for performance improvement, and efficiently retrieving information for audiences and analysts. While sports video reasoning is clearly important, it is under-explored due to challenges and the lack of datasets.
@@
In this paper, we tackle sports VideoQA, focusing on reasoning over sports videos for question answering. This is distinct from general VideoQA, which centers on general video understanding, mainly encompassing descriptive and temporal aspects [15, 41, 39, 46], as depicted in Fig. 2. Compared to general VideoQA, sports VideoQA is more challenging because the questions may involve particular teams and professional actions, and it requires a fine-grained understanding of actions and intra/inter-team relation modeling. As pointed out in [33]: 1) coarse-grained global video understanding is insufficient for sports analysis where the actions are rapid and hard to differentiate; 2) the models pre-trained on coarse-grained datasets are not always beneficial to fine-grained tasks. In the meantime, although significant strides have been made in the fine-grained recognition of individual actions [33, 23, 32], these datasets may fall short in effectively evaluating the ability for temporal and causal reasoning over actions performed by individuals or teams in sports scenarios. A large-scale dataset containing professional sports scenarios with detailed annotations of sports-relevant questions is thus required for explicit and comprehensive reasoning for sports VideoQA.
To address this gap, we introduce a novel dataset called Sports-QA, the first video QA dataset specifically focusing on sports activities. Leveraging sports videos and professional action labels from MultiSports [18] and FineGym [33], we construct our dataset. Both sources provide high-quality sports videos with well-defined action labels, allowing us to annotate essential attributes for each crucial action instance. Using these annotations, we generate QA pairs based on pre-defined templates covering aspects such as description, chronology, causality, and counterfactual conditions. The Sports-QA dataset is the first of its kind, comprising approximately 94K QA pairs, encompassing multiple sports and various question types for sports VideoQA. Table 1 provides a comparison between Sports-QA and several widely-used VideoQA datasets.
@@
MSRVTT-QA [41] General 10K 244K Auto OE
MSVD-QA [41] 2K 51K Auto OE
NExT-QA [39] 5.4K 52K Man OE&MC
TGIF-QA [15] 72K 165K Auto&Man OE&MC
MovieQA [35] Movie 408 15K Man MC
Social-IQ [47] Social 1K 8K Man MC
DramaQA [3] Drama 24K 16K Auto&Man MC
@@ 2 Related Work
VideoQA Datasets. The development of VideoQA has been greatly facilitated by the emergence of various datasets, such as TGIF-QA [15], MSVD-QA, and MSRVTT-QA [41], DramaQA [3], NExT-QA [39], NExT-OOD [50]. TGIF-QA, for instance, offers a comprehensive set of four sub-tasks designed for temporal reasoning in general videos and is widely recognized in the VideoQA community. MSVD-QA and MSRVTT-QA are open-ended datasets constructed from existing video captions, while DramaQA focuses on understanding drama stories with hierarchical QAs and character-centered video annotation. NExT-QA, introduced for describing and explaining temporal actions, provides both multi-choice and open-ended questions through manual annotation. However, existing datasets predominantly center around general video understanding in daily scenarios or involve coarse-grained action/event reasoning. In contrast, our contribution lies in fine-grained and professional analysis within sports scenarios. We present Sports-QA, a dataset that addresses aspects of description, chronology, causality, and counterfactual conditions for multiple sports of diverse characteristics.
@@
VideoQA Methods. VideoQA poses a significant challenge as it necessitates models to grasp both spatial and temporal information from videos to answer questions. Various deep models have been developed, approaching this task from different perspectives [8, 14, 24, 1, 10, 42, 44, 25, 20]. For instance, the deep heterogeneous graph alignment network by Jiang et al. [14] addresses VideoQA by simultaneously aligning intra/inter-modality information. Another approach involves a multimodal attention model proposed by Fan et al. [8], where heterogeneous memory learns global context from visual features, and question memory captures the complex semantics of questions. IGV, introduced by Li et al. [24], grounds question-critical scenes in videos by considering causal relations that remain invariant to complement contents. However, a limitation in these approaches is the lack of consideration for the fact that different questions may require temporal dependencies of specific scales. In response to this challenge, we propose the Auto-Focus Transformer, designed to automatically focus on a specific temporal scale based on the question for VideoQA.
@@
Sports Video Understanding. Sports video understanding has drawn increasing attention in recent years [45, 19, 17, 49, 26, 38]. Researchers have made great efforts in various tasks such as sports action recognition [33], multi-person action detection [18], and action quality assessment [34]. Meanwhile, numerous sports datasets are constructed [27, 18, 9, 5]. For example, AQA-7 [27] is constructed for professional action quality assessment. Seven types of actions are included in this dataset and all the action instances are associated with quality scores. MultiSports [18] is proposed for the spatial-temporal detection of professional sports actions. FineGym [33] focuses on fine-grained activity localization that requires an accurate understanding of the atomic level of a gymnastic action. In this work, we construct Sports-QA, the first dataset for sports VideoQA, bridging the gap between the visual modality and the textual modality in sports analysis.
@@ 3.1 Data Source
Regarding the collection of sports videos, we consider the following aspects: 1) The visual quality of the video data should be high enough to conduct fine-grained video understanding, such as video resolution and frame rate. 2) Instead of applying to a single type of sports, we expect a VideoQA dataset involving multiple sports. After a deep survey of the works on sports video understanding, we find that the MultiSports [18] and FineGym [33] datasets are highly suitable for our purpose. The details of MultiSports and FineGym are as follows.
MultiSports [18] is a dataset for the temporal localization of sports actions, encompassing four sports (i.e., basketball, football, volleyball, and aerobic gymnastics) and 66 fine-grained action categories [^3] Project homepage: https://deeperaction.github.io/datasets/multisports.html. The action categories in MultiSports are professional terms (such as “volleyball spike”, “football tackle”, and “basketball defensive rebound”) instead of common and atomic actions like “run” and “stand”. For each action instance, the dataset provides the bounding boxes of a player from the starting frame of an action to the ending frame, forming the action spatial-temporal tube. The statistics of MultiSports are shown in Table 2.
FineGym [33] is a dataset designed to elevate the field of action recognition by addressing the limitations observed in existing techniques [^4] Project homepage: https://sdolivia.github.io/FineGym. Developed to surpass current benchmarks, it offers a unique combination of richness, quality, and diversity in its content. This dataset is constructed on gymnasium videos, providing a realistic and varied environment for action recognition studies. What sets FineGym apart from other datasets is its meticulous temporal annotation at both action and sub-action levels, featuring a three-level semantic hierarchy. This hierarchical structure allows for a more nuanced understanding of activities, enabling researchers to explore and analyze actions in finer detail. As an illustrative example, an event is annotated as a sequence of elementary sub-actions Importantly, each sub-action within these sets is further annotated with finely defined class labels, contributing to a higher level of granularity in action recognition. Fig. 2 shows the action hierarchy of the MultiSports and FineGym datasets.
@@ 3.2 Question-Answer Pair Generation
In most existing works, QA pairs are generated using two approaches: 1) Automatically generating questions and answers based on video captions using a set of predefined question templates [15, 41, 43]. This method is efficient, but it may introduce obvious grammatical errors or lose crucial information during the conversion of captions into questions and answers. 2) Manually annotating questions and answers through crowdsourcing [39, 46, 11]. Although this approach produces QA pairs of high quality in terms of accuracy and expression, manual annotation is time-consuming and expensive. In this work, we aim to ensure the quality of the textual data while considering annotation costs. We achieve this by generating QA pairs using pre-defined templates based on existing labels and newly-labeled attributes of actions. Our approach guarantees grammatically error-free QA pairs that capture crucial information in videos, thanks to the careful design of sophisticated templates. Furthermore, as ball games and gymnastics have distinct characteristics, we generate the QA pairs in different ways as described as follows.
@@ 3.2.1 QA Pair Generation for Ball Games
Attribute Definition. MultiSports consists of 45 action categories from ball games, with some being crucial for professional sports statistics or quantifying the performance of players/teams. For instance, the “2-Point Field Goal Percentage” in a basketball match necessitates a fine-grained understanding of all “2-point shot” actions, while the “saving” actions in football directly reflect the performance of goalkeepers. In this work, we focus on 28 crucial actions, listed in Table LABEL:acts. Considering both practical applications and research purposes, we define five attributes for each of the crucial action instances as follows.
@@
2. Outcome: This attribute assesses whether the result of the action meets the expectation and is defined as a binary label: successful or failed. For example, if a “2-point shot” in basketball scores a goal, it is annotated as successful; otherwise, it is labeled as failed.
3. Cause of Outcome: This attribute indicates the cause of failure for offensive actions, such as a “2-point shot” in basketball, “shoot” in football, and “spike” in volleyball, or the cause of success for defensive actions, like “block” in football and “save” in volleyball. Specifically, the cause of failure for offensive actions reflects player shortcomings. For example, the failure of an offensive “long pass” in football could be attributed to defensive interception, bad pass, or bad catch. Identifying the precise cause helps teams identify their weaknesses and improve performance. Similarly, the cause of success for defensive actions highlights player strengths. For instance, the success of a defensive “save” in volleyball might be due to offensive out of bounds, offensive blocked by net, or actually catch the ball, with the last case directly reflecting defensive performance. We do not consider the cause of success for offensive actions or the cause of failure for defensive actions, as they cannot be attributed to different cases. For example, the failure of “football saving” can only be explained by the goalkeeper missing the ball. Therefore, these causes are not worth discussing. We define various cause options for applicable action categories and ask annotators to choose from them. Note that this attribute is applicable to certain action categories, and the causes vary for different action categories.
@@
From Attributes to Question-Answer Pairs. In particular, the questions are queries about the attributes, and their answers can be obtained directly from our annotations or through logical reasoning and statistical analysis over the annotations. For instance, for a question asking about the cause of a specific action, the answer can be easily retrieved from its attribute of cause; For a question asking whether a specific action is successful, we can check its outcome attribute. We guarantee the correctness of the answers as we have checked the annotations of attributes as mentioned previously. These answers include responses like “yes/no”, numerical values, sports names (such as “volleyball”), action categories in MultiSports and FineGym (such as “block”), and short phrases describing the reasons (such as “out of bounds”). See Fig. 3 for more concrete examples of the attributes and the corresponding generated QA pairs.

### v5

@@ Abstract
Reasoning over sports videos for question answering is an important task with numerous applications, such as player training and information retrieval. However, this task has not been explored due to the lack of relevant datasets and the challenging nature it presents. Most datasets for video question answering (VideoQA) focus mainly on general and coarse-grained understanding of daily-life videos, which is not applicable to sports scenarios requiring professional action understanding and fine-grained motion analysis. In this paper, we introduce the first dataset, named Sports-QA, specifically designed for the sports VideoQA task. The Sports-QA dataset includes various types of questions, such as descriptions, chronologies, causalities, and counterfactual conditions, covering multiple sports. Furthermore, to address the characteristics of the sports VideoQA task, we propose a new Auto-Focus Transformer (AFT) capable of automatically focusing on particular scales of temporal information for question answering. We conduct extensive experiments on Sports-QA, including baseline studies and the evaluation of different methods. The results demonstrate that our AFT achieves state-of-the-art performance [^1] Sports-QA is available at: https://github.com/HopLee6/Sports-QA.
@@ 1 Introduction
Sports video analysis has been attracting increasing attention in recent years [yuan2021spatio, li2021groupformer, koshkina2021contrastive, zhu2022fencenet, martin2020fine, wang2022shuttlenet, 6516867]. While research progress has been made on tasks such as sports action recognition [li2021multisports, zhu2022fencenet, rasmussen2022compressing], reasoning over sports videos for question answering has not been explored. As humans, we can not only recognize the actions of the players in sports videos but also understand the effects of the players’ actions, explain why a team loses the score, and imagine what would happen under counterfactual situations. Our impressive capabilities in reasoning allow us to answer complex questions related to sports videos, which is crucial in applications such as obtaining crucial statistics in matches for player/team performance qualification, analyzing players’ actions and team strategies for performance improvement, and efficiently retrieving information for audiences and analysts. While sports video reasoning is clearly important, it is under-explored due to challenges and the lack of datasets.
@@
In this paper, we tackle sports VideoQA, focusing on reasoning over sports videos for question answering. This is distinct from general VideoQA, which centers on general video understanding, mainly encompassing descriptive and temporal aspects [jang2017tgif, xu2017video, xiao2021next, yu2019activitynet], as depicted in Fig. 3. Compared to general VideoQA, sports VideoQA is more challenging because the questions may involve particular teams and professional actions, and it requires a fine-grained understanding of actions and intra/inter-team relation modeling. As pointed out in [shao2020finegym]: 1) coarse-grained global video understanding is insufficient for sports analysis where the actions are rapid and hard to differentiate; 2) the models pre-trained on coarse-grained datasets are not always beneficial to fine-grained tasks. In the meantime, although significant strides have been made in the fine-grained recognition of individual actions [shao2020finegym, liu2022fineaction, sun2017taichi], these datasets may fall short in effectively evaluating the ability for temporal and causal reasoning over actions performed by individuals or teams in sports scenarios. A large-scale dataset containing professional sports scenarios with detailed annotations of sports-relevant questions is thus required for explicit and comprehensive reasoning for sports VideoQA.
To address this gap, we introduce a novel dataset called Sports-QA, the first video QA dataset specifically focusing on sports activities. Leveraging sports videos and professional action labels from MultiSports [li2021multisports] and FineGym [shao2020finegym], we construct our dataset. Both sources provide high-quality sports videos with well-defined action labels, allowing us to annotate essential attributes for each crucial action instance. Using these annotations, we generate QA pairs based on pre-defined templates covering aspects such as description, chronology, causality, and counterfactual conditions. The Sports-QA dataset is the first of its kind, comprising approximately 94K QA pairs, encompassing multiple sports and various question types for sports VideoQA. Table 1 provides a comparison between Sports-QA and several widely-used VideoQA datasets.
@@
MSRVTT-QA [xu2017video] General 10K 244K Auto OE
MSVD-QA [xu2017video] 2K 51K Auto OE
NExT-QA [xiao2021next] 5.4K 52K Man OE&MC
TGIF-QA [jang2017tgif] 72K 165K Auto&Man OE&MC
SportQA [xia2024sportqa] Sports 0 70K Auto&Man MC
SPORTU-text [xia2024sportu] Sports 0 0.9K Man MC
SPORTU-video [xia2024sportu] Sports 1.7K 12K Man MC
MovieQA [tapaswi2016movieqa] Movie 408 15K Man MC
Social-IQ [zadeh2019social] Social 1K 8K Man MC
DramaQA [choi2021dramaqa] Drama 24K 16K Auto&Man MC
@@ 2.1 VideoQA Datasets
The development of VideoQA has been greatly facilitated by the emergence of various datasets, such as TGIF-QA [jang2017tgif], MSVD-QA, and MSRVTT-QA [xu2017video], DramaQA [choi2021dramaqa], NExT-QA [xiao2021next], NExT-OOD [10107423]. TGIF-QA, for instance, offers a comprehensive set of four sub-tasks designed for temporal reasoning in general videos and is widely recognized in the VideoQA community. MSVD-QA and MSRVTT-QA are open-ended datasets constructed from existing video captions, while DramaQA focuses on understanding drama stories with hierarchical QAs and character-centered video annotation. NExT-QA, introduced for describing and explaining temporal actions, provides both multi-choice and open-ended questions through manual annotation. However, existing datasets center around general video understanding in daily scenarios or involve coarse-grained action/event reasoning. Our contribution lies in fine-grained and professional analysis within sports scenarios. We present Sports-QA, a dataset that addresses aspects of description, chronology, causality, and counterfactual conditions for multiple sports of diverse characteristics.
2.2 VideoQA Methods
@@
VideoQA poses a significant challenge as it necessitates models to grasp both spatial and temporal information from videos to answer questions. Various deep models have been developed, approaching this task from different perspectives [fan2019heterogeneous, jiang2020reasoning, li2022invariant, antol2015vqa, gao2018motion, 10172254, 9770842, 10214041, 10146482]. For instance, the deep heterogeneous graph alignment network by Jiang et al. [jiang2020reasoning] addresses VideoQA by simultaneously aligning intra/inter-modality information. Another approach involves a multimodal attention model proposed by Fan et al. [fan2019heterogeneous], where heterogeneous memory learns global context from visual features, and question memory captures the complex semantics of questions. IGV, introduced by Li et al. [li2022invariant], grounds question-critical scenes in videos by considering causal relations that remain invariant to complement contents. However, a limitation in these approaches is the lack of consideration for the fact that different questions may require temporal dependencies of specific scales. In response to this challenge, we propose the Auto-Focus Transformer, designed to automatically focus on a specific temporal scale based on the question.
2.3 Sports Video/Text Understanding
@@
Sports video/text understanding has drawn increasing attention in recent years [yuan2021spatio, li2021groupformer, koshkina2021contrastive, zhu2022fencenet, martin2020fine, wang2022shuttlenet, yang2024sports, wiseman2017challenges]. Researchers have made great efforts in various tasks such as sports action recognition [shao2020finegym], multi-person action detection [li2021multisports], and action quality assessment [tang2020uncertainty]. Meanwhile, numerous sports datasets are constructed [parmar2019action, li2021multisports, giancola2018soccernet, deliege2021soccernet]. AQA-7 [parmar2019action] is constructed for professional action quality assessment. Seven types of actions are included in this dataset and all the action instances are associated with quality scores. A large-scale dataset of basketball game descriptions paired with detailed statistics is introduced by [wiseman2017challenges], enabling fully data-driven modeling approaches. MultiSports [li2021multisports] is proposed for the spatial-temporal detection of professional sports actions. FineGym [shao2020finegym] focuses on fine-grained activity localization that requires an accurate understanding of the atomic level of a gymnastic action. Sports Intelligence [yang2024sports] focuses on comprehensive LLM evaluation in sports-related NLP tasks. While it covers diverse textual scenarios, it does not address the multimodal or temporal complexities of sports video QA. In contrast, our dataset is the first to specifically benchmark LLMs in the context of professional sports video understanding, offering a new challenge for multimodal reasoning.
2.4 Sports VideoQA
SportQA [xia2024sportqa] presents over 70,000 multiple-choice questions to evaluate sports knowledge in LLMs. However, it is purely text-based and lacks any video content. In contrast, our Sports-QA is a video-centric QA benchmark, targeting the unique temporal and visual reasoning challenges present in sports. SPORTU [xia2024sportu] introduces both text- and video-based QA. However, SPORTU-video consists of slow-motion clips with generic QA focused on rule application and basic event recognition. Our Sports-QA differs in three key ways: 1) It uses high-tempo professional sports broadcasts, which are more fine-grained and realistic. 2) It introduces structured question types aligned with video understanding (e.g., descriptive, chronological, causal, and counterfactual reasoning). 3) It targets higher-level semantic reasoning grounded in temporal dynamics, not just static recognition.
@@ 3.1 Data Source
Regarding the collection of sports videos, we consider the following aspects: 1) The visual quality of the video data should be high enough to conduct fine-grained video understanding, such as video resolution and frame rate. 2) Instead of applying to a single type of sports, we expect a VideoQA dataset involving multiple sports. After a deep survey of the works on sports video understanding, we find that the MultiSports [li2021multisports] and FineGym [shao2020finegym] datasets are highly suitable for our purpose. The details of MultiSports and FineGym are as follows.
MultiSports [li2021multisports] is a dataset for the temporal localization of sports actions, encompassing four sports (i.e., basketball, football, volleyball, and aerobic gymnastics) and 66 fine-grained action categories [^4] Project homepage: https://deeperaction.github.io/datasets/multisports.html. The action categories in MultiSports are professional terms (such as “volleyball spike”, “football tackle”, and “basketball defensive rebound”) instead of common and atomic actions like “run” and “stand”. For each action instance, the dataset provides the bounding boxes of a player from the starting frame of an action to the ending frame, forming the action spatial-temporal tube. The statistics of MultiSports are shown in Table 2.
FineGym [shao2020finegym] is a dataset designed to elevate the field of action recognition by addressing the limitations observed in existing techniques [^5] Project homepage: https://sdolivia.github.io/FineGym. Developed to surpass current benchmarks, it offers a unique combination of richness, quality, and diversity in its content. This dataset is constructed on gymnasium videos, providing a realistic and varied environment for action recognition studies. What sets FineGym apart from other datasets is its meticulous temporal annotation at both action and sub-action levels, featuring a three-level semantic hierarchy. This hierarchical structure allows for a more nuanced understanding of activities, enabling researchers to explore and analyze actions in finer detail. As an illustrative example, an event is annotated as a sequence of elementary sub-actions Importantly, each sub-action within these sets is further annotated with finely defined class labels, contributing to a higher level of granularity in action recognition. Fig. 2 shows the action hierarchy of the MultiSports and FineGym datasets.
@@ 3.2 Question-Answer Pair Generation
In most existing works, QA pairs are generated using two approaches: 1) Automatically generating questions and answers based on video captions using a set of predefined question templates [jang2017tgif, xu2017video, yang2021just]. This method is efficient, but it may introduce obvious grammatical errors or lose crucial information during the conversion of captions into questions and answers. 2) Manually annotating questions and answers through crowdsourcing [xiao2021next, yu2019activitynet, garcia2020knowit]. Although this approach produces QA pairs of high quality in terms of accuracy and expression, manual annotation is time-consuming and expensive. In this work, we aim to ensure the quality of the textual data while considering annotation costs. We achieve this by generating QA pairs using pre-defined templates based on existing labels and newly-labeled attributes of actions. Our approach guarantees grammatically error-free QA pairs that capture crucial information in videos, thanks to the careful design of sophisticated templates. Furthermore, as ball games and gymnastics have distinct characteristics, we generate the QA pairs in different ways as described as follows.
@@ 3.2.1 QA Pair Generation for Ball Games
Attribute Definition. MultiSports consists of 45 action categories from ball games, with some being crucial for professional sports statistics or quantifying the performance of players/teams. For instance, the “2-Point Field Goal Percentage” in a basketball match necessitates a fine-grained understanding of all “2-point shot” actions, while the “saving” actions in football directly reflect the performance of goalkeepers. In this work, we focus on 28 crucial actions, listed in Table 11. Considering both practical applications and research purposes, we define five attributes for each of the crucial action instances as follows.
@@
2. Outcome: This attribute assesses if the action meets the expectation and is defined as a binary label: successful or failed. For example, if a “2-point shot” in basketball scores a goal, it is annotated as successful; otherwise, it is labeled as failed.
3. Cause of Outcome:
This attribute indicates the cause of failure for offensive actions, such as a “2-point shot” in basketball, “shoot” in football, and “spike” in volleyball, or the cause of success for defensive actions, like “block” in football and “save” in volleyball. Specifically, the cause of failure for offensive actions reflects player shortcomings. For example, the failure of an offensive “long pass” in football could be attributed to defensive interception, bad pass, or bad catch. Identifying the precise cause helps teams identify their weaknesses and improve performance. Similarly, the cause of success for defensive actions highlights player strengths. For instance, the success of a defensive “save” in volleyball might be due to offensive out of bounds, offensive blocked by net, or actually catch the ball, with the last case directly reflecting defensive performance. We do not consider the cause of success for offensive actions or the cause of failure for defensive actions, as they cannot be attributed to different cases. For example, the failure of “football saving” can only be explained by the goalkeeper missing the ball. Therefore, these causes are not worth discussing. We define various cause options for applicable action categories and ask annotators to choose from them. Note that this attribute is applicable to certain action categories, and the causes vary for different action categories.
@@ 3. Cause of Outcome:
From Attributes to Question-Answer Pairs. In particular, the questions are queries about the attributes, and their answers can be obtained directly from our annotations or through logical reasoning and statistical analysis over the annotations. For instance, for a question asking about the cause of a specific action, the answer can be easily retrieved from its attribute of cause; For a question asking whether a specific action is successful, we can check its outcome attribute. We guarantee the correctness of the answers as we have checked the annotations of attributes as mentioned previously. These answers include responses like “yes/no”, numerical values, sports names (such as “volleyball”), action categories in MultiSports and FineGym (such as “block”), and short phrases describing the reasons (such as “out of bounds”). See Fig. 3 for more concrete examples of the attributes and the generated QA pairs.